dx/dt=x-y

dy/dt=x+3y


1
Expert's answer
2021-04-29T16:28:33-0400

We can express the system in the form "X'=AX" , where:


"X=\begin{bmatrix}
 x \\
 y
\end{bmatrix},\quad X'=\begin{bmatrix}
 x' \\
 y'
\end{bmatrix},\quad A=\begin{bmatrix}
 1 & -1 \\
 1 & 3
\end{bmatrix}"

We found a normal-shaped Jordan matrix similar to the original:


"|A-\lambda I|=\begin{vmatrix}
 1-\lambda & -1 \\
 1 & 3-\lambda
\end{vmatrix}=(\lambda -2)^2"

Therefore, its eigenvalue is "\lambda=2" with algebraic multiplicity "2".

We find the eigenvectors of eigenvalue "\lambda=2":


"v=(-1,1)"

We know that for repeated real eigenvalues "\lambda" with multiplicity "2" and eigenvector "v" , the general solution takes the form:


"X=c_1e^{\lambda t}v+c_2e^{\lambda t}(tv+u)"

where "u" is a solution to "(A-\lambda I)u=v". We can see that:


"u=(1,0)"

Finally, the solution is:


"X=c_1e^{2t}\begin{bmatrix}
 -1 \\
 1
\end{bmatrix}+c_2e^{2t}\left( t+\begin{bmatrix}
 -1 \\
 1
\end{bmatrix}+\begin{bmatrix}
 1 \\
 0
\end{bmatrix}\right)."